# Sequential Attention-based Sampling for Histopathological Analysis

### v3

@@ 4.2 Baselines and Metrics
We evaluate and compare the performance of our model against two types of baselines. First, we evaluate a range of different sota methods that analyze all patches in the WSI at high resolution, including MaxPooling, ABMIL [ilse2018abmil], CLAM [lu2021clam], TransMIL [shao2021transmil], DTFD [zhang2022dtfd], and ACMIL [zhang2023acmil]. We also include comparisons with our hierarchical attention-based (HAFED) model (Section˜3.2) with the full-resolution WSI. Yet, these methods serve only as an upper bound on achievable performance.
For a fair comparison, we also evaluate models that sample only a fraction of the WSI. Specifically, we compare our approach with RLogist [zhao2022rlogist] when trained to sample exactly $10\%$ and $20\%$ of patches at high resolution, under identical experimental settings. We also compare with a non-RL-based Zoom-MIL [zoom_mil] model that samples exactly k = 80, 160, or 300 patches at low resolution (5x $\rightarrow$ 20x). The choice of k values aligns with the patch statistics of the dataset: WSIs contain $\sim$800 low-resolution patches on average; thus, $k=80$ and $k=160$ correspond to $10\%$ and $20\%$ of the available patches, on average, providing a fair comparison with the RLogist-0.1/SASHA-0.1 and RLogist-0.2/SASHA-0.2 settings, respectively. $k=300$ follows the default configuration used in the original Zoom-MIL paper [zoom_mil]. This setup allows us to systematically analyze the effect of varying observation budgets on model performance. In addition, we also compare against ACMIL and DTFD under the same observation budgets, with random sampling of high-resolution patches (Table˜2); details in Appendix A.3. Models were trained with the official codebase provided by the respective studies, except for RLogist. The RLogist codebase is insufficient for replication; therefore, we report the performance metrics as is from the original study [zhao2022rlogist].
@@ 4.4 Ablation Study and Limited Sampling
We performed a systematic ablation study removing each component of the model in turn – or replacing it with a naive variant – and evaluating it with the CAMELYON16 dataset. These included: a) replacing the WSI-pretrained ViT with a ResNet50, used commonly for feature extraction [zheng2024cvpr, zhao2022rlogist], b) replacing ViT with the CONCH medical image-encoder[lu2023conch], pretrained on 1.17M image caption pairs, c) replacing multi-branch attention with single branch in the Classifer [ilse2018abmil], d) replacing the targeted state update policy with a global state update [zhao2022rlogist], e) removing the TSU and updating the state only of the sampled (local, high-resolution) patch, and f) and selecting an action based on a random policy (see “Variant”s in Table˜2). In every case, we observed sharp drops in accuracy, ranging from 7-45%, relative to the baseline SASHA model; similar drops occurred in AUROC and F1 scores Table˜2. The strongest impact occurred when changing the RL policy (Table˜2, f). The next strongest impact occurred when changing the feature extractor to ResNet-50 (a), but this impact was mitigated when using the CONCH encoder (b). Intermediate levels of impact occurred when changing the classifier to single-branch attention (c) or altering the state update method (d-e). Additional ablation experiments are described in Appendix A.3. These results indicate that every component of our model was critical for achieving high performance, comparable to sota models.
@@ 4.6 Effect of observation budget: Calibration
We have shown that SASHA achieves near-sota performance at a fraction of the inference time, even with sampling 10% or 20% of the patches at high resolution. Is there any advantage to sampling more patches, other than achieving a marginal improvement accuracy, especially when the inference time increases almost linearly with the proportion of high-resolution patches sampled. Here, we show empirically that sampling a higher fraction of patches yields better calibrated predictions. With well calibrated models, the model’s confidence in its prediction (e.g., predictive uncertainty) is a useful indicator for when referral to a specialist is warranted [Dj_and_jim]. We quantified the expected calibration error (ECE) [guo2017calibration] as a measure of the degree of calibration of the models. The ECE decreased systematically at the proportion sampled patches increased from 10% (SASHA-0.1) to 20% (SASHA-0.2) to 100% (HAFED). This pattern occurred possibly because providing the model more timesteps to sample the WSI in each RL episode enabled the model to recover from sub-optimal sampling actions in the early timesteps, and to sample the relevant portions of the WSI in the later timesteps. Surprisingly, ECE with SASHA-0.2 was lower than that of even sota attention-based models like ACMIL and DTFD (Fig.˜3h). In other words, while a shorter observation budget encouraged efficient learning, a longer budget produced better calibrated predictions.
@@ Acknowledgements
This research was supported by a Department of Science and Technology SwarnaJayanti Fellowship, a Google Research Grant, and a Pratiksha Trust Intramural Grant (all to DS), Ministry of Education Scholarships (to TG and NM), an Institute of Eminence (IoE) Grant (to IISc), a Walmart Centre for Tech Excellence at IISc, the Indo–French Centre for the Promotion of Advanced Research Grant (CEFIPRA; Project 7102–1), Kotak IISc FinTech Grant, a DST–SERB Core Research Grant (CRG/2021/008330), and a Pratiksha Trust Young Investigator Award (all to GT). We are very grateful to Karm Patel, Mainak Biswas, Aditi Saxena, and Tej Pratap Yadav for help with model development and simulations, and Raj Jain for help with the original draft of the manuscript. The authors declare no competing interests.
Sequential Attention-based Sampling for Histopathological Analysis
@@ A.1 Dataset details
The CAMELYON16 dataset comprises 399 whole-slide images (WSIs) of lymph node sections. After pre-processing and segmentation, we generated non-overlapping patches of size $256\times 256$ ($\times 3$ RGB channels). On average, this resulted in, on average, $\sim$800 patches per WSI at a scanning resolution of $5\times$, and $\sim$12,500 patches at a high resolution of $20\times$. We used the official test set for evaluation [bejnordi2017diagnostic]. The training set was further divided to create a validation set (10%) while maintaining a similar proportion of normal and tumor slides in each subset.
The TCGA-NSCLC dataset comprises 1,054 whole-slide images (WSIs) obtained from the TCGA-NSCLC repository under the TCGA-LUSC and TCGA-LUAD projects. Ten low-quality slides without magnification information were discarded, and three slides with errors were ignored. We followed the same test split as in DSMIL [dsmil2021]. The test set comprised a total of 213 slides: 104 from TCGA-LUAD and 109 from TCGA-LUSC. Following preprocessing, we obtained, on average, $\sim$212 patches per WSI at $5\times$ magnification and $\sim$3,393 patches at $20\times$ magnification. The training set was further divided to create a validation (20%) set while maintaining a similar proportion of labels in each subset.
@@
The CamelyonPlus dataset[ling2025camelyonplus], comprises 1,347 whole‐slide images (WSIs) derived from the curated CAMELYON16 and CAMELYON17 collections, after removal of low-quality slides and re-labeling. It involves a four-class classification. We created training, validation, and test splits to ensure class balance across all four labels. A summary of the splits and labels can be found in Table˜4.
@@
The BRACS (BReAst Carcinoma Subtyping) dataset comprises 547 labeled WSIs obtained from the public repository [brancati2022bracs]. This dataset involves a 7-class classification task, characterizing each tumor as one of three benign (BT), two atypical (AT), or two malignant (MT) sub-types. We ignored the “normal” subtype in the benign (BT) class, and each set of subtypes within BT, AT, and MT was grouped into its respective class. Additionally, a small number of slides (7/503) that lacked all relevant levels of magnification were excluded. After preprocessing, we use the official dataset split across train, validation, and test split with details outlined in Table˜5.
@@ A.2.1 Patch feature distillation with HAFED
The first step of SASHA involves feature extraction and aggregation of high-resolution patch features, which specifies the RL agent’s state $S(t)$. We use a Vision Transformer (ViT) pretrained on histopathological images to extract features from each zoomed in, high-resolution patch [kang2023ssl]. This ViT was used as is without any additional fine tuning, and its output – a 384 dimensional embedding per patch – was provided as input to a hierarchical attention-based feature aggregator and distiller (HAFED).
@@
HAFED is trained with an annealed learning rate schedule starting at 4 $\times 10^{-4}$, with the AdamW optimizer and a weight decay of 1 $\times 10^{-4}$. The HAFED module is trained end-to-end using the following loss components: (1) a cross-entropy loss between the final predicted label and the ground truth label, (2) a cross-entropy loss applied independently to the predictions from each attention branch in second stage and ground truth label, and (3) a dissimilarity loss that encourages diversity among the attention scores of the different branches in the second stage. This dissimilarity loss encourages each branch to focus on distinct tumor features, thereby improving the model’s ability to capture complementary diagnostic features[zhang2023acmil].
@@ A.2.3 Actor and Critic
The final component of SASHA is a reinforcement learning (RL) agent responsible for sampling, sequentially, the most relevant patches for high-resolution analysis. This selection is formulated as a sequential decision-making problem, where the agent learns to identify diagnostically informative regions within each slide. We adopt the Actor-Critic framework and train the agent using the Proximal Policy Optimization (PPO) algorithm, a widely used on-policy method known for its stability and effectiveness in complex RL tasks [schulman2017ppo].
@@
The Critic network also mirrors the backbone of actor network i.e architecture of Layers 1–3 in Table˜6, with an additional final layer that takes the aggregated state representation ($\alpha\cdot S_{t}$), and predicts the reward-to-go or expected return given the current state $S_{t}$. The reward-to-go is used in calculating the generalized advantage estimate (GAE, [gae]). During training the actual reward-to-go from the given state is known, and the critic network is trained by minimizing MSE loss between the predicted and actual value.
@@ A.3.2 Alternative Models under an Observation Budget
We compared the performance of SASHA with competing sota models like ACMIL [zhang2023acmil] and DTFD [zhang2022dtfd] under the same observation budget of either 10% or 20%. That is, we predicted slide labels by having these sota models examine only a limited fraction of patches in each WSI at high resolution, matching the fraction of the WSI sampled by SASHA.
@@ A.3.4 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section˜A.2.3). In contrast, a previous competing model (RLogist [zhao2022rlogist]) employed a terminal reward strategy, providing feedback only at the end of each episode.
@@
Reward Schedule Accuracy AUC F1
Top-k attention score
patches $\cap$ RL actions
Avg attention
score of RL actions
@@ A.4 Computation of WSI Compressibility, Inference time, and Calibration Error
Expected Calibration Error (ECE): Expected Calibration Error (ECE) is a metric used to quantify how well a neural network’s predicted probabilities align with the true probabilities (or observed frequencies) of its predictions [guo2017calibration]. It measures how reliable a model’s confidence levels are in its predictions. A perfectly calibrated model would have an ECE of 0, indicating that the model’s confidence scores accurately reflect the true likelihood of its predictions being correct.
To compute ECE, we first divide the predicted confidence scores (i.e., the maximum softmax probabilities) into $M=10$ equally spaced bins. For each bin, we calculate two quantities: i) Accuracy: the proportion of correctly classified samples in the bin; and ii) Confidence: the average of the predicted probabilities in the bin. These metrics are then used to compute the calibration error as the absolute difference between its accuracy and confidence for each bin, summed across bins [guo2017calibration].

### v4

@@ 4.2 Baselines and Metrics
We evaluate and compare the performance of our model against two types of baselines. First, we evaluate a range of different sota methods that analyze all patches in the WSI at high resolution, including MaxPooling, ABMIL [15], CLAM [22], TransMIL [32], DTFD [38], and ACMIL [39]. We also include comparisons with our hierarchical attention-based (HAFED) model (Section˜3.2) with the full-resolution WSI. Yet, these methods serve only as an upper bound on achievable performance.
For a fair comparison, we also evaluate models that sample only a fraction of the WSI. Specifically, we compare our approach with RLogist [40] when trained to sample exactly $10\%$ and $20\%$ of patches at high resolution, under identical experimental settings. We also compare with a non-RL-based Zoom-MIL [36] model that samples exactly k = 80, 160, or 300 patches at low resolution (5x $\rightarrow$ 20x). The choice of k values aligns with the patch statistics of the dataset: WSIs contain $\sim$800 low-resolution patches on average; thus, $k=80$ and $k=160$ correspond to $10\%$ and $20\%$ of the available patches, on average, providing a fair comparison with the RLogist-0.1/SASHA-0.1 and RLogist-0.2/SASHA-0.2 settings, respectively. $k=300$ follows the default configuration used in the original Zoom-MIL paper [36]. This setup allows us to systematically analyze the effect of varying observation budgets on model performance. In addition, we also compare against ACMIL and DTFD under the same observation budgets, with random sampling of high-resolution patches (Table˜2); details in Appendix A.3. Models were trained with the official codebase provided by the respective studies, except for RLogist. The RLogist codebase is insufficient for replication; therefore, we report the performance metrics as is from the original study [40].
@@ 4.4 Ablation Study and Limited Sampling
We performed a systematic ablation study removing each component of the model in turn – or replacing it with a naive variant – and evaluating it with the CAMELYON16 dataset. These included: a) replacing the WSI-pretrained ViT with a ResNet50, used commonly for feature extraction [41, 40], b) replacing ViT with the CONCH medical image-encoder[24], pretrained on 1.17M image caption pairs, c) replacing multi-branch attention with single branch in the Classifer [15], d) replacing the targeted state update policy with a global state update [40], e) removing the TSU and updating the state only of the sampled (local, high-resolution) patch, and f) and selecting an action based on a random policy (see “Variant”s in Table˜2). In every case, we observed sharp drops in accuracy, ranging from 7-45%, relative to the baseline SASHA model; similar drops occurred in AUROC and F1 scores Table˜2. The strongest impact occurred when changing the RL policy (Table˜2, f). The next strongest impact occurred when changing the feature extractor to ResNet-50 (a), but this impact was mitigated when using the CONCH encoder (b). Intermediate levels of impact occurred when changing the classifier to single-branch attention (c) or altering the state update method (d-e). Additional ablation experiments are described in Appendix A.3. These results indicate that every component of our model was critical for achieving high performance, comparable to sota models.
@@ 4.6 Effect of observation budget: Calibration
We have shown that SASHA achieves near-sota performance at a fraction of the inference time, even with sampling 10% or 20% of the patches at high resolution. Is there any advantage to sampling more patches, other than achieving a marginal improvement accuracy, especially when the inference time increases almost linearly with the proportion of high-resolution patches sampled. Here, we show empirically that sampling a higher fraction of patches yields better calibrated predictions. With well calibrated models, the model’s confidence in its prediction (e.g., predictive uncertainty) is a useful indicator for when referral to a specialist is warranted [8]. We quantified the expected calibration error (ECE) [11] as a measure of the degree of calibration of the models. The ECE decreased systematically at the proportion sampled patches increased from 10% (SASHA-0.1) to 20% (SASHA-0.2) to 100% (HAFED). This pattern occurred possibly because providing the model more timesteps to sample the WSI in each RL episode enabled the model to recover from sub-optimal sampling actions in the early timesteps, and to sample the relevant portions of the WSI in the later timesteps. Surprisingly, ECE with SASHA-0.2 was lower than that of even sota attention-based models like ACMIL and DTFD (Fig.˜3h). In other words, while a shorter observation budget encouraged efficient learning, a longer budget produced better calibrated predictions.
@@ Acknowledgements
This research was supported by a Department of Science and Technology SwarnaJayanti Fellowship, a Google Research Grant, a Pratiksha Trust Intramural Grant and a Gore Subraya Bhat Chair Associate Professorship in Digital Health (all to DS), Ministry of Education Scholarships (to TG and NM), an Institute of Eminence (IoE) Grant (to IISc), a Walmart Centre for Tech Excellence at IISc, the Indo–French Centre for the Promotion of Advanced Research Grant (CEFIPRA; Project 7102–1), Kotak IISc FinTech Grant, a DST–SERB Core Research Grant (CRG/2021/008330), and a Pratiksha Trust Young Investigator Award (all to GT). We are very grateful to Karm Patel, Mainak Biswas, Aditi Saxena, and Tej Pratap Yadav for help with model development and simulations, and Raj Jain for help with the original draft of the manuscript. The authors declare no competing interests.
@@ A.1 Dataset details
The CAMELYON16 dataset comprises 399 whole-slide images (WSIs) of lymph node sections. After pre-processing and segmentation, we generated non-overlapping patches of size $256\times 256$ ($\times 3$ RGB channels). On average, this resulted in, on average, $\sim$800 patches per WSI at a scanning resolution of $5\times$, and $\sim$12,500 patches at a high resolution of $20\times$. We used the official test set for evaluation [3]. The training set was further divided to create a validation set (10%) while maintaining a similar proportion of normal and tumor slides in each subset.
The TCGA-NSCLC dataset comprises 1,054 whole-slide images (WSIs) obtained from the TCGA-NSCLC repository under the TCGA-LUSC and TCGA-LUAD projects. Ten low-quality slides without magnification information were discarded, and three slides with errors were ignored. We followed the same test split as in DSMIL [18]. The test set comprised a total of 213 slides: 104 from TCGA-LUAD and 109 from TCGA-LUSC. Following preprocessing, we obtained, on average, $\sim$212 patches per WSI at $5\times$ magnification and $\sim$3,393 patches at $20\times$ magnification. The training set was further divided to create a validation (20%) set while maintaining a similar proportion of labels in each subset.
@@
The CamelyonPlus dataset[21], comprises 1,347 whole‐slide images (WSIs) derived from the curated CAMELYON16 and CAMELYON17 collections, after removal of low-quality slides and re-labeling. It involves a four-class classification. We created training, validation, and test splits to ensure class balance across all four labels. A summary of the splits and labels can be found in Table˜4.
@@
The BRACS (BReAst Carcinoma Subtyping) dataset comprises 547 labeled WSIs obtained from the public repository [4]. This dataset involves a 7-class classification task, characterizing each tumor as one of three benign (BT), two atypical (AT), or two malignant (MT) sub-types. We ignored the “normal” subtype in the benign (BT) class, and each set of subtypes within BT, AT, and MT was grouped into its respective class. Additionally, a small number of slides (7/503) that lacked all relevant levels of magnification were excluded. After preprocessing, we use the official dataset split across train, validation, and test split with details outlined in Table˜5.
@@ A.2.1 Patch feature distillation with HAFED
The first step of SASHA involves feature extraction and aggregation of high-resolution patch features, which specifies the RL agent’s state $S(t)$. We use a Vision Transformer (ViT) pretrained on histopathological images to extract features from each zoomed in, high-resolution patch [17]. This ViT was used as is without any additional fine tuning, and its output – a 384 dimensional embedding per patch – was provided as input to a hierarchical attention-based feature aggregator and distiller (HAFED).
@@
HAFED is trained with an annealed learning rate schedule starting at 4 $\times 10^{-4}$, with the AdamW optimizer and a weight decay of 1 $\times 10^{-4}$. The HAFED module is trained end-to-end using the following loss components: (1) a cross-entropy loss between the final predicted label and the ground truth label, (2) a cross-entropy loss applied independently to the predictions from each attention branch in second stage and ground truth label, and (3) a dissimilarity loss that encourages diversity among the attention scores of the different branches in the second stage. This dissimilarity loss encourages each branch to focus on distinct tumor features, thereby improving the model’s ability to capture complementary diagnostic features[39].
@@ A.2.3 Actor and Critic
The final component of SASHA is a reinforcement learning (RL) agent responsible for sampling, sequentially, the most relevant patches for high-resolution analysis. This selection is formulated as a sequential decision-making problem, where the agent learns to identify diagnostically informative regions within each slide. We adopt the Actor-Critic framework and train the agent using the Proximal Policy Optimization (PPO) algorithm, a widely used on-policy method known for its stability and effectiveness in complex RL tasks [31].
@@
The Critic network also mirrors the backbone of actor network i.e architecture of Layers 1–3 in Table˜6, with an additional final layer that takes the aggregated state representation ($\alpha\cdot S_{t}$), and predicts the reward-to-go or expected return given the current state $S_{t}$. The reward-to-go is used in calculating the generalized advantage estimate (GAE, [30]). During training the actual reward-to-go from the given state is known, and the critic network is trained by minimizing MSE loss between the predicted and actual value.
@@ A.3.2 Alternative Models under an Observation Budget
We compared the performance of SASHA with competing sota models like ACMIL [39] and DTFD [38] under the same observation budget of either 10% or 20%. That is, we predicted slide labels by having these sota models examine only a limited fraction of patches in each WSI at high resolution, matching the fraction of the WSI sampled by SASHA.
@@ A.3.4 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section˜A.2.3). In contrast, a previous competing model (RLogist [40]) employed a terminal reward strategy, providing feedback only at the end of each episode.
@@
Reward Schedule Accuracy AUC F1 Top-k attention score patches $\cap$ RL actions Avg attention score of RL actions
@@ A.4 Computation of WSI Compressibility, Inference time, and Calibration Error
Expected Calibration Error (ECE): Expected Calibration Error (ECE) is a metric used to quantify how well a neural network’s predicted probabilities align with the true probabilities (or observed frequencies) of its predictions [11]. It measures how reliable a model’s confidence levels are in its predictions. A perfectly calibrated model would have an ECE of 0, indicating that the model’s confidence scores accurately reflect the true likelihood of its predictions being correct.
To compute ECE, we first divide the predicted confidence scores (i.e., the maximum softmax probabilities) into $M=10$ equally spaced bins. For each bin, we calculate two quantities: i) Accuracy: the proportion of correctly classified samples in the bin; and ii) Confidence: the average of the predicted probabilities in the bin. These metrics are then used to compute the calibration error as the absolute difference between its accuracy and confidence for each bin, summed across bins [11].